OpenStudy (anonymous):

HOW DO I FIND THE ANSWER TO THIS (1,3)AND(3,13) M=

OpenStudy (nenadmatematika):

m=(13-3)/(3-1)=10/2=5

OpenStudy (anonymous):

OK SO HOW DID YOU FIGURE THAT HUN? I DON'T UNDERSTAND LINEAR FUNCTIONS

OpenStudy (nenadmatematika):

draw axis x and y and put this coordinates, draw the line between through the points and m is the slope of the line which you can get by formula I gave you...try to work it out, it's not hard, it's basics

OpenStudy (anonymous):

m = slope = rise/run = the change in y divided by the change in x = dy/dx = (y1-y2)/(x1-x2)

OpenStudy (anonymous):

OK LET ME TRY THIS NEXT PROBLEM AND SEE IF I GET IT

OpenStudy (anonymous):

(X,Y)=(-1,3)____=M(___)_____ =M

OpenStudy (anonymous):

OH BOY I DON'T GET IT LOL

OpenStudy (anonymous):

take off your caps lock so it doesn't seem like you are yelling at everyone

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

it's okay aha..

OpenStudy (nenadmatematika):

hahaha :D

OpenStudy (anonymous):

I forget about it to be honest

OpenStudy (anonymous):

I know lol. it's totally cool so let's help you with your problem now that you're in a talking voice :) ahha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I still don't understand this